Manuela has a total of 40 dvds of movies and tv shows. The number of movies is 4 less than 3 times the number of tv shows. Write and solve a system of equations to find the numbers of movies and tv show that she has on dvd.
Answer by JulietG(1812)

Let M = movies; let T = TV shows
M + T = 40 (Manuela has a total of 40 dvds of movies and tv shows.)
M = 3T - 4 (4 less than 3 times the number of tv shows.)
Take the known value of M from the bottom equation and put it into the top.
M + T = 40
(3T-4) + T = 40
4T - 4 =40
Add 4 to each side
4T = 44
Divide each side by 4
T = 11
If there are 11 TV shows, and 40 in total, then there are 29 movies
29 = 3(11) - 4
29 = 29